OpenStudy (anonymous):

Suppose that x=6 and y=8. If y is inversely related to x, what it y when x=3? 4 5 16 18

OpenStudy (anonymous):

What *is not What it, I know how to do direct variation but i get confused with inverse relations, please if someone could explain this to me that would be awesome!

OpenStudy (phi):

How do you do *direct variation* ?

OpenStudy (anonymous):

Well direct variation has the form y=kx where k is not equal to zero

OpenStudy (phi):

inverse variation is almost the same, except inverse means *divide* by x y = k/x

OpenStudy (anonymous):

Right how do i apply to this problem?

OpenStudy (phi):

if you they give you a pair of numbers (and they do) x=6 and y=8 replace x with 6 and y with 8 in y = k/x can you do that ?

OpenStudy (anonymous):

8=k/6, okay where do i go from there? do plug in 3 for k?

OpenStudy (phi):

next step is solve for k. to do that you multiply both sides by 6 like this 8*6 = 6* k/6 the right side has 6*k/6 which can be written as k*6/6 (change the order of multiplying) and then notice 6/6 is 1 you get 8*6= k or k= 48 once you found the "magic number" k , you use it by putting in the number in y = k/x in other words, your inverse variation equation is y= 48/x now if you are given an x (a new number) put that number in for x in the equation and simplify to find y what is y when x is 3 ?

OpenStudy (anonymous):

oh so then y=48/3 = 16!! so y is 16 when x is 3, that makes sense got it!! now i can solve the rest of these prob thanks!!!

OpenStudy (phi):

these problems take 3 steps: 1) write down the equation 2) solve for k using the info they give you 3) use k in the equation, and use it to answer the question

OpenStudy (anonymous):

Thanks a lot @phi !! ur a lot better at explaining then teacher or textbook!

OpenStudy (phi):

yw